A Level Mathematics 9709 Paper 53, May/June 2026

Question paper and mark scheme50 marks6 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9709/53 · May/June 2026

Download PDF
1

A fair 8-sided dice with faces labelled 1, 2, 3, 4, 5, 6, 7, 8 is thrown repeatedly.

a

Find the probability that a 5 will be obtained for the first time on the 7th throw.

[1]
b

Find the probability that a 5 will be obtained for the first time before the 10th throw.

[2]
c

Find the probability that a 5 will be obtained at least once but no more than three times in 12 randomly chosen throws.

[3]
2

The random variable X takes the values −2,0,1,3. It is given that P(X=x)=k(x+1)2, where k is a constant.

a

Draw up the probability distribution table for X, giving the probabilities as numerical fractions not involving k.

[3]
b

Find the value of Var⁡(X).

[3]
3

The times taken, in minutes, by 150 students to complete a particular computer game are summarised in the table.

time taken (t minutes)0⩽t<1010⩽t<1515⩽t<2020⩽t<2525⩽t<3535⩽t<60
frequency42440462610
a

Draw a histogram to represent this information.

[4]
b

Calculate an estimate for the mean of these times.

[3]
c

In which class interval does the 40th percentile of the times lie?

[1]
4

Ebere has a coin which is biased so that the probability of obtaining a head is 14. She also has two bags containing coloured discs. Bag A contains 5 red discs and 3 blue discs. Bag B contains 4 red discs and 5 green discs. Ebere tosses her coin. If she obtains a head, she chooses two discs from bag A, at random, one at a time and without replacement. If she obtains a tail, she chooses two discs from bag B, at random, one at a time and without replacement.

a

Draw a fully labelled tree diagram to represent this information.

[2]
b

Find the probability that Ebere obtains two discs of the same colour.

[3]
c

Find the probability that Ebere obtained a head with her coin given that at least one of the discs is red.

[4]
5
a

The lengths of Indian cobra snakes are normally distributed with mean 125 cm and standard deviation 8 cm. Find the probability that a randomly chosen Indian cobra has a length between 105 cm and 135 cm.

[3]
b

The masses of Indian cobras are normally distributed with mean 2.5 kg and standard deviation σ kg. It is known that 81% of Indian cobras have a mass less than 2.7 kg. Find the value of σ.

[3]
c

A random sample of 80 Indian cobras is taken. Use an approximation to find the probability that fewer than 60 of these 80 Indian cobras have a mass less than 2.7 kg.

[5]
6
a

Find the number of different arrangements of the 10 letters in the word POSSESSION in which the four Ss are all next to each other and the two Os are not next to each other.

[3]
b

Find the number of different arrangements of the 10 letters in the word POSSESSION in which the P is at the beginning, the N is at the end and there are exactly four letters between the two Os.

[3]
c

Five letters are selected from the 10 letters in the word POSSESSION. Find the number of different selections in which the five letters include at least two Ss and at least one O.

[4]

Mark scheme

9709/53 · May/June 2026

Download PDF

1a · 1 mark

  • B1

    [(78)6×18]=0.0561,1176492097152

1b · 2 marks

  • M1

    1−(78)9 or full addition of terms up to power 8

  • A1

    = 0.699

1c · 3 marks

  • M1

    One term of the form 12Cx(p)x(1−p)12−x, 0<p<1, x≠0 or 12.

  • A1

    Correct un-simplified expression.

  • B1

    = 0.746

2a · 3 marks

  • B1

    [22k=1] k=122

  • B1

    Table with correct values of x and at least one non-zero probability correct or FT their k, 0<k<1.

  • B1

    4 correct probabilities linked with the correct outcomes.

2b · 3 marks

  • M1

    [E(X)=]−2×k+[0+]1×4k+3×16k

  • M1

    Var(X)=1×(−2)2+[1×02+]4×12+16×3222−(their E(X))2

  • A1

    = 1.74

3a · 4 marks

  • M1

    At least 4 frequency densities calculated.

  • A1

    All bar heights correct on graph, NOT FT, daylight rule, using their suitable linear scale with at least 3 values indicated. At least 50% of the grid must be used for each axis.

  • B1

    Bar ends at 10, 15, 20, 25, 35, 60. All 6 bars drawn, with at least 3 values indicated, 0⩽horizontal scale⩽60. At least 50% of the grid must be used for each axis.

  • B1

    Axes labelled frequency density (fd) time (t) and minutes, OE, or an appropriate title. (Axes may be reversed).

3b · 3 marks

  • B1

    Midpoints: 5 12.5 17.5 22.5 30 47.5

  • M1

    Correct un-simplified mean formula with their 6 midpoints (not upper bound, lower bound, upper limit, lower limit, class width, frequency density, frequency and must be within class), accept un-simplified. Condone 1 value error on numerator.

  • A1

    =22.1,22115,2210150

3c · 1 mark

  • B1

    15⩽t<20

4a · 2 marks

  • B1

    Fully correct structure labelled tree diagram. Probabilities for coin correct 14,34.

  • B1

    Probabilities for Bag A and Bag B correct (irrespective of correct coin toss being present). First draw 58,38,49,59. Second draw 47,37,57,27,38,58,48,48.

4b · 3 marks

  • M1

    a4×b8×c7 or 8 or d4×e9×f8 or 9 seen or a4×b56 or 64 or c4×d72 or 81 seen.

  • M1

    3 correct scenarios added – maybe identified on tree diagram or by seeing P(RR)+P(BB)+P(GG).

  • A1

    151336,0.449

4c · 4 marks

  • M1

    P(H∩1R or 2R)=14×58×47+14×58×37+14×38×57

  • M1

    P(1R or 2R)=1−P(BB)−P(GG)

  • M1

    Correct fraction calculation seen – accept un-simplified, or FT their clearly identified P(H∩1R or 2R) and P(1R or 2R) used to form conditional probability.

  • A1

    75257,0.292

5a · 3 marks

  • M1

    Use of ±standardisation formula once with 125, 8 and either 105 or 135. No continuity correction. No σ2,σ.

  • M1

    Calculating the appropriate probability area, must be a probability. Values given here come from tables and may be slightly different if using values from a calculator. Leading to their final answer, expect > 0.5.

  • A1

    = 0.888

5b · 3 marks

  • B1

    0.8775⩽z<0.8785, −0.8785<z⩽−0.8775 seen.

  • M1

    2.7 and 2.5 substituted in ± standardisation formula equated to a z-value, (±0.2σ) equated to a z-value scores M1 BOD. No continuity correction, not σ2,σ.

  • A1

    [σ=]100439,0.228

5c · 5 marks

  • B1

    64.8 and 12.312 or 12.31 seen, allow un-simplified. May be in standardisation formula. (σ=12.312,3.5088,99525…to at least 4SF implies correct variance).

  • M1

    Substituting their 64.8 and their 12.312 into ±standardisation formula, any number for 59.5, allow σ2,σ.

  • M1

    Use continuity correct 59.5 or 60.5 in their standardisation formula.

  • M1

    Appropriate area Φ, from final process, must be a probability. May be implied by a sketch of the required probability areas. Note: correct final answer implies this M1. Expect final answer < 0.5.

  • A1

    = 0.0655

6a · 3 marks

  • M1

    Method 1: a−6!,a>720. Or a−4!×5×6,a>720, a must be an integer. Method 2: 5!×m,m>1.

  • M1

    Method 1: 7!c!−(k!×n,n=1 or 2),k=5 or 6 and c=1 or 2. Method 2: n× 6C2 or n× 6C4 or n×6×52 or n×6P22!. Condone n× 6P2 or n× 6P4 or n×6×5, n must be an integer.

  • A1

    = 1800

6b · 3 marks

  • B1

    Method 1: 6!4!×b,b⩾1,b an integer. Method 2: (6P44! or  6C4)×d,d⩾1,d an integer.

  • M1

    Method 1: c×3,c⩾6,c an integer. Method 2: e×3!,e⩾6,e an integer.

  • A1

    = 90

6c · 4 marks

  • M1

    One identified outcome 4Ck×n,k=1,2,3 and n⩾1.

  • M1

    3 outcomes of the form 4Ck×n,k=1,2 and n⩾1 un-simplified.

  • M1

    5 correct scenarios added (values do not need to be correct) may be identified by un-simplified expressions.

  • A1

    [Number of selections = ] 16

Questions

  1. 1
    Question 16 marks
    Mark scheme ›
  2. 2
    Question 26 marks
    Mark scheme ›
  3. 3
    Question 38 marks
    Mark scheme ›
  4. 4
    Question 49 marks
    Mark scheme ›
  5. 5
    Question 511 marks
    Mark scheme ›
  6. 6
    Question 610 marks
    Mark scheme ›

More 9709 papers

Other papers from May/June 2026

Solve it on ExamPalStart